# Direct-CP: Directed Collaborative Perception for Connected and Autonomous Vehicles via Proactive Attention

### v2

@@ III-C Direction-weighted detection loss
Given the final fused feature map $\mathcal{F}_{i}^{out}$, the detection decoder $\Phi_{dec}(\cdot)$ generates class and regression outputs following [10]. Each output location $\Phi_{dec}(\mathcal{F}_{i}^{out})\in\mathbb{R}^{H\times W\times 7}$ corresponds to a rotated box described by a 7-tuple $(c,x,y,h,w,\cos\alpha,\sin\alpha)$, representing class confidence, position, size, and angle. To evaluate the discrepancy between the collaborative 3D detection results and the ground truth, the commonly used detection loss $\mathcal{L}_{det}$ [21] combines focal loss, object offset loss, and object size loss. However, this loss does not fully capture the importance of specific directions in our directed CP scenario. Therefore, we introduce a novel loss function, direction-weighted detection loss (DWLoss), to quantify the divergence in designated directions. DWLoss is calculated by dividing the 3D detection results into $N_{dir}$ subsets and computing the detection loss $\{\mathcal{L}_{det}^{i}\}_{i=1}^{N_{dir}}$ for each subset with varying sum weights, represented as follows:
@@
where $\sigma$ is a constant weight-control factor. Eq. 8 ensures setting lower weights to non-critical directions by weight factor $\sigma$, aiming to jointly optimize the CP performance in interested directions and the remaining directions. The choice of $\sigma$ is crucial: setting it too high may obscure the importance of interested directions, while setting it too low (an extreme case is 0), can neglect the accuracy in non-critical directions during training, potentially degrading perception more than single-vehicle perception. Ablation studies in Section IV will offer helpful guidance for determining an effective $\sigma$.
@@
Figure 4: Visualization comparison on V2X-Sim 2.0 dataset. Green and red boxes denote ground truth and predictions, respectively. While Where2comm improves global perception over lower-bound, it shows limitations in certain directions. Direct-CP enhances perception in ego’s interested directions (marked as 1, with right arrow showing ego’s movement).
@@ IV-A Experimental setup
Dataset and baselines. Our experimental evaluations are conducted on the V2X-Sim 2.0 Dataset [22], an extensive simulated dataset generated using the CARLA simulator [23]. This dataset comprises 10,000 frames of 3D LiDAR point clouds along with 501,000 annotated 3D bounding boxes. We configure the perception range to be $64m\times 64m$, and the 3D points are discretized into a BEV map of dimensions $(252,100,64)$. We establish baseline comparisons, including When2com [5], V2VNet [8], and Where2comm [10]. To make the perception gain clearer, we set the single-vehicle perception method as the lower-bound baseline.
Implementation details. We implement our Direct-CP using PyTorch. The direction control module features a fully connected layer with dimensions of $100\times 252$ for both input and output, complemented by a sigmoid layer to match the BEV feature dimensions, incorporating pose matrices and direction masks at a spatial resolution of $(100,252)$. Our detection module utilizes the LiDAR-based 3D object detection framework PointPillar [24]. We set the training batch size at 6 and the maximum epochs at 60. The 360-degree space is divided into 4 directions: [0, 90], [90, 180], [180, 270], and [270, 360], corresponding to left front, right front, right back, and left back, with interest weights of [0.9, 0.9, 0.1, 0.1], respectively. The default DWLoss weight factor $\sigma$ is 1.0 and the default communication budget (defined in Eq. 3) is 0.2. The setup for our experiments includes 2 Intel(R) Xeon(R) Silver 4410Y CPUs (2.0GHz), 4 NVIDIA RTX A5000 GPUs, and 512GB DDR4 RAM.
@@ IV-B Quantitative results
Figure 5: Attention weight visualization on neighboring CAVs. Where2comm distributes attention equally for omnidirectional perception, while Direct-CP focuses on features relevant to ego’s interested directions.
Evaluation of Direct-CP. We evaluate Direct-CP against baselines in the overall CP performance (AP@IoU=0.5/0.7) and in specific directions (AP@PD-IoU=0.5/0.7, interested directions are denoted with *). As shown in Table II, Direct-CP uses direction-aware selective attention to reallocate communication resources, slightly outperforming the state-of-the-art Where2comm in terms of overall AP@IoU. For PD-IoU, Where2comm optimizes CP omnidirectionally, showing similar AP@PD-IoU across all directions, while Direct-CP focuses on preferred directions, achieving 18.2% higher AP@PD-IoU=0.5 and 19.8% higher AP@PD-IoU=0.7 than Where2comm in these directions. These results demonstrate that Direct-CP enables an ego vehicle to flexibly adjust view focus and improve CP performance in the desired directions.
@@
When2comm [5] 34.97 33.56 19.36 49.96 53.56 24.81 24.59 8.75 40.42 38.70
V2VNet [8] 57.49 53.62 28.01 60.36 67.35 42.61 41.05 19.54 36.59 48.22
Where2comm [10] 51.29 59.38 48.83 56.27 79.59 45.86 44.52 37.89 48.83 64.96
Direct-CP (Ours) 65.84+28.4% 65.48+10.3% 37.21 60.55 81.17+2.0% 55.76+21.6% 53.20+19.5% 28.62 49.98 66.57+2.5%
@@
Communication efficiency. Moreover, we investigate how varying communication budgets affects CP performance as shown in Fig. 3 with budgets ranging from 0.01 to 0.25. Notably, below a budget of 0.1, both Direct-CP and Where2comm experience a significant drop in AP@IoU=0.7 and AP@PD-IoU=0.7 for interested directions [0, 180]. Despite this, Direct-CP slightly outperforms Where2comm overall and significantly improves perception in interested directions. At a further reduced budget of 0.01, both methods perform equally, suffering major perception degradation likely due to ultra-sparse feature maps impeding model convergence. Overall, these results highlight Direct-CP’s efficiency under constrained communication resources.
Abalation studies. To investigate the influence of the weight factor $\sigma$ on the performance of Direct-CP, we conduct an ablation study, varying $\sigma$ from 0 to 2.0. When $\sigma$ is below 1.0, we observe a reduction in collaborative detection accuracy, particularly in less critical directions. Notably, AP@PD-IoU=0.7 for the sector [270, 360] declines to 0.03, markedly deteriorating below the lower-bound threshold. Conversely, when $\sigma$ exceeds 1.5, there is a discernible decrease in detection accuracy for both the areas of interest and the overall system. Based on these observations, a good range for $\sigma$ is between 1.0 and 1.5, which balances directed perception performance with satisfactory overall accuracy.
@@ IV-C Qualitative results
Visualization of collaborative 3D detection results. As shown in Fig. 4, we display Direct-CP’s collaborative detection results alongside baselines on the V2X-Sim 2.0 dataset. While Where2comm substantially improves global perception over the lower-bound, it underperforms in certain local directions, occasionally not exceeding single-vehicle outcomes, likely due to limited communication budgets and scattered focus. Conversely, our Direct-CP effectively redirects attention from less critical to key areas, significantly boosting local directional perception.
Visualization of ego CAV’s attention weights. As depicted in Fig. 5, we further compare the attention weights of ego CAV assigned to neighboring CAVs’ feature maps $W_{i,j}^{DSA}$ (defined in Eq. 6) in two methods. With limited communication budgets, both methods query sparse features. For Where2comm, the attention weights are more uniformly assigned to other CAVs to enhance 360-degree CP performance. In contrast, our proposed Direct-CP attends to features that are more crucial to the ego vehicle’s interested directions, informed by other CAVs’ pose information and the ego’s directional mask, shifting great attention from CAV 2 and 4 to CAV 1 and 3 to improve directed CP performance.
@@
In this paper, we have introduced Direct-CP, a novel CP system for ego vehicles to enhance perception in patronized directions. We have developed RSU-aided direction masking by integrating RSU’s traffic detection with ego vehicle’s interests to identify key directions. We have also designed a proactive direction-aware attention mechanism to collect sparse feature maps from multiple vehicles under limited communication budgets, improving local directional perception. Additionally, we have created a direction-weighted detection loss to align perception outputs with ground truth. Extensive experiments demonstrate that Direct-CP achieves directed performance gains under constrained communication resources and outperforms baselines in efficiency. Looking forward, our direction-aware framework opens new possibilities for adaptive perception, and future work could explore incorporating sophisticated traffic indicators and extending to extreme weather conditions and complex urban environments.

### v4

@@ III-C Direction-weighted detection loss
Given the final fused feature map $\mathcal{F}_{i}^{out}$, the detection decoder $\Phi_{dec}(\cdot)$ generates class and regression outputs following [hu2022wherecomm]. Each output location $\Phi_{dec}(\mathcal{F}_{i}^{out})\in\mathbb{R}^{H\times W\times 7}$ corresponds to a rotated box described by a 7-tuple $(c,x,y,h,w,\cos\alpha,\sin\alpha)$, representing class confidence, position, size, and angle. The angle is parameterized using both sine and cosine components to ensure continuous optimization and avoid the periodicity issue at $\pm\pi$. To evaluate the discrepancy between the collaborative 3D detection results and the ground truth, the commonly used detection loss $\mathcal{L}_{det}$ [zhou2019objectspoints] combines focal loss, object offset loss, and object size loss. However, this loss does not fully capture the importance of specific directions in our directed CP scenario. Therefore, we introduce a novel loss function, direction-weighted detection loss (DWLoss), to quantify the divergence in designated directions. DWLoss is calculated by dividing the 3D detection results into $N_{dir}$ subsets and computing the detection loss $\{\mathcal{L}_{det}^{i}\}_{i=1}^{N_{dir}}$ for each subset with varying sum weights, represented as follows:
@@
where $\sigma$ is a constant weight-control factor. Eq. 8 ensures that lower weights are assigned to non-critical directions by weight factor $\sigma$, aiming to jointly optimize the CP performance in interested directions and the remaining directions. The choice of $\sigma$ is crucial: too high may obscure the importance of interested directions, while too low (an extreme case is 0) can neglect the accuracy in non-critical directions during training, potentially degrading perception more than single-vehicle perception. Ablation studies in Section IV will offer helpful guidance for determining an effective $\sigma$.
@@
Figure 4: Visualization comparison on V2X-Sim 2.0 dataset. Green and red boxes denote ground truth and predictions, respectively. While Where2comm improves global perception over lower-bound, it shows limitations in certain directions. Directed-CP enhances perception in ego’s interested directions (marked as 1, with right arrow showing ego CAV’s movement).
@@ IV-A Experimental setup
Dataset and baselines. Our experimental evaluations are conducted on the V2X-Sim 2.0 Dataset [li2022v2xsimmultiagentcollaborativeperception], an extensive simulated dataset generated using the CARLA simulator [pmlr-v78-dosovitskiy17a]. This dataset comprises 10,000 frames of 3D LiDAR point clouds along with 501,000 annotated 3D bounding boxes. We configure the perception range to be $64m\times 64m$, and the 3D points are discretized into a BEV map of dimensions $(252,100,64)$. We establish baseline comparisons, including When2com [Liu_2020_CVPR], V2VNet [10.1007/978-3-030-58536-5_36], and Where2comm [hu2022wherecomm]. To make the perception gain clearer, we set the single-vehicle perception method as the lower-bound baseline.
Implementation details. We implement our Directed-CP using PyTorch. The direction control module features a fully connected layer with dimensions of $100\times 252$ for both input and output, complemented by a sigmoid layer to match the BEV feature dimensions, incorporating pose matrices and direction masks at a spatial resolution of $(100,252)$. Our detection module utilizes the LiDAR-based 3D object detection framework PointPillar [Lang_2019_CVPR]. We set the training batch size at 6 and the maximum epochs at 60. The 360-degree space is divided into 4 directions: [0, 90], [90, 180], [180, 270], and [270, 360], corresponding to left front, right front, right back, and left back, with interest weights of [0.9, 0.9, 0.1, 0.1], respectively. The default DWLoss weight factor $\sigma$ is 1.0 and the default communication budget (defined in Eq. 3) is 0.2. The setup for our experiments includes 2 Intel(R) Xeon(R) Silver 4410Y CPUs (2.0GHz), 4 NVIDIA RTX A5000 GPUs, and 512GB DDR4 RAM.
@@ IV-B Quantitative results
Figure 5: Attention weight visualization on neighboring CAVs. Where2comm distributes attention equally for omnidirectional perception, while Directed-CP focuses on features relevant to ego’s interested directions.
Evaluation of Directed-CP. We evaluate Directed-CP against baselines in the overall CP performance (AP@IoU=0.5/0.7) and in specific directions (AP@PD-IoU=0.5/0.7, interested directions are denoted with *). As shown in Table II, Directed-CP uses direction-aware selective attention to reallocate communication resources, slightly outperforming the state-of-the-art Where2comm in terms of overall AP@IoU. For PD-IoU, Where2comm optimizes CP omnidirectionally, showing similar AP@PD-IoU across all directions, while Directed-CP focuses on preferred directions, achieving 18.2% higher AP@PD-IoU=0.5 and 19.8% higher AP@PD-IoU=0.7 than Where2comm in these directions. These results demonstrate that Directed-CP enables an ego CAV to flexibly adjust view focus and improve CP performance in the desired directions.
@@
When2comm [Liu_2020_CVPR] 34.97 33.56 19.36 49.96 53.56 24.81 24.59 8.75 40.42 38.70
V2VNet [10.1007/978-3-030-58536-5_36] 57.49 53.62 28.01 60.36 67.35 42.61 41.05 19.54 36.59 48.22
Where2comm [hu2022wherecomm] 51.29 59.38 48.83 56.27 79.59 45.86 44.52 37.89 48.83 64.96
Directed-CP (Ours) 65.84+28.4% 65.48+10.3% 37.21 60.55 81.17+2.0% 55.76+21.6% 53.20+19.5% 28.62 49.98 66.57+2.5%
@@
Communication efficiency. We investigate how varying communication budgets affects CP performance as shown in Fig. 3 with budgets ranging from 0.01 to 0.25. Notably, below a budget of 0.1, both Directed-CP and Where2comm experience a significant drop in AP@IoU=0.7 and AP@PD-IoU=0.7 for interested directions [0, 180]. Despite this, Directed-CP slightly outperforms Where2comm overall and significantly improves perception in interested directions. At a further reduced budget of 0.01, both methods perform equally, suffering major perception degradation likely due to ultra-sparse feature maps impeding model convergence. Overall, these results highlight Directed-CP’s efficiency under constrained communication resources.
Abalation studies. To investigate the influence of the weight factor $\sigma$ on the performance of Directed-CP, we conduct an ablation study, varying $\sigma$ from 0 to 2.0. When $\sigma$ is below 1.0, we observe a reduction in collaborative detection accuracy, particularly in less critical directions. Notably, AP@PD-IoU=0.7 for the sector [270, 360] declines to 0.03, markedly deteriorating below the lower-bound threshold. Conversely, when $\sigma$ exceeds 1.5, there is a discernible decrease in detection accuracy for both the areas of interest and the overall system. Based on these observations, a good range for $\sigma$ is between 1.0 and 1.5, which balances directed perception performance with satisfactory overall accuracy.
@@ IV-C Qualitative results
Visualization of collaborative 3D detection results. As shown in Fig. 4, we display Directed-CP’s collaborative detection results alongside baselines on the V2X-Sim 2.0 dataset. While Where2comm substantially improves global perception over the lower-bound, it underperforms in certain local directions, occasionally not exceeding single-vehicle outcomes, likely due to limited communication budgets and scattered focus. Conversely, our Directed-CP effectively redirects attention from less critical to key areas, significantly boosting local directional perception.
Visualization of ego CAVs’ attention weights. As depicted in Fig. 5, we further compare the attention weights that the ego CAV assigned to neighboring CAVs’ feature maps $W_{i,j}^{DSA}$ (defined in Eq. 6) in two methods. With limited communication budgets, both methods query sparse features. For Where2comm, the attention weights are more uniformly assigned to other CAVs to enhance 360-degree CP. In contrast, our proposed Directed-CP pays attention to features that are more crucial to the ego CAV’s interested directions, informed by other CAVs’ pose information and the ego CAV’s directional mask, shifting great attention from CAV 2 and 4 to CAV 1 and 3 to improve directed CP.
@@
In this paper, we have introduced Directed-CP, a novel CP system for ego CAVs to enhance perception in patronized directions. We have developed RSU-aided direction masking by integrating RSU’s traffic detection with ego CAV’s interests to identify key directions. We have also designed a proactive direction-aware attention to collect sparse feature maps from surrounding CAVs under communication budgets to improve directional perception. Additionally, we have created a direction-weighted detection loss to align perception outputs with ground truth. Experiments demonstrate that Directed-CP achieves controllable and directed perception gains under constrained communication and outperforms baselines in efficiency. Looking forward, our direction-aware framework opens new possibilities for adaptive perception, and future work could explore incorporating sophisticated traffic indicators and extending to more complex urban environments under extreme weather conditions.